Find the maximum value of the function f(x) = x + x'2 − x'3
for x ≥ 0
1
Expert's answer
2019-09-10T09:37:34-0400

Find the maximum value of the function "f(x)=x+x^2-x^3" for "x\u22650".

Solution

Find the critical points of the function:


"f'(x)=1+2x-3x^2=0;"


"-3x^2+2x+1=0."

We solve the resulting equation using discriminants


"D=2^2-4\cdot(-3)\cdot1=16;"

"x_1=\frac{-2+\sqrt{16}} {-3\cdot2}=-\frac1 3;"

"x_2=\frac{-2-\sqrt{16}} {-3\cdot2}=1."

Option "-\frac1 3" is not suitable, because "x\u22650".

Now we define the intervals of increasing ("f'(x)>0") and decreasing ("f'(x)<0") functions.by the interval method.


We get that the point "x=1" is the maximum.

"f(1)=1+1^2-1^3=1."

Answer

The maximum value of the function is 1.